How to calculate pH of solution containing 50 mL of 1M HCl and 100mL of 0.25 M HCl?

Grade:12th pass

1 Answers

Arun
6 years ago
Resulting molarity = 
M * 150 = 1 *50 + 0.25 * 100
M * 150 = 50 + 25
M = 75/150
M = 0.5
 
Now I hope you can solve this and get pH.